# Using LLMs in Generating Design Rationale for Software Architecture Decisions

### v1

@@ Abstract.
Design Rationale (DR) for software architecture decisions refers to the reasoning underlying architectural choices, which provides valuable insights into the different phases of the architecting process throughout software development. However, in practice, DR is often inadequately documented due to a lack of motivation and effort from developers. With the recent advancements in Large Language Models (LLMs), their capabilities in text comprehension, reasoning, and generation may enable the generation and recovery of DR for architecture decisions. In this study, we evaluated the performance of LLMs in generating DR for architecture decisions. First, we collected 50 Stack Overflow (SO) posts, 25 GitHub issues, and 25 GitHub discussions related to architecture decisions to construct a dataset of 100 architecture-related problems. Then, we selected five LLMs to generate DR for the architecture decisions with three prompting strategies, including zero-shot, chain of thought (CoT), and LLM-based agents. With the DR provided by human experts as ground truth, the Precision of LLM-generated DR with the three prompting strategies ranges from 0.267 to 0.278, Recall from 0.627 to 0.715, and F1-score from 0.351 to 0.389. Additionally, 64.45% to 69.42% of the arguments of DR not mentioned by human experts are also helpful, 4.12% to 4.87% of the arguments have uncertain correctness, and 1.59% to 3.24% of the arguments are potentially misleading. Based on the results, we further discussed the pros and cons of the three prompting strategies and the strengths and limitations of the DR generated by LLMs.
@@ 1. Introduction
The software architecture of a system consists of the structures required to understand the system, including its software elements, their relationships and properties (Bass et al., 2021). When making architecture decisions, software architects need to consider multiple factors like the application domain, architectural styles and patterns, Commercial Off-The-Shelf (COTS) components, and other infrastructure choices required to meet system requirements (Jansen and Bosch, 2005). Design Rationale (DR) explains the reasoning behind architecture decisions, encapsulating the architecture knowledge and thought processes that support the resulting design (Tang et al., 2006). A comprehensive DR can support various development activities, such as change impact analysis or a major redesign. With documented DR, architecture decisions can be more easily revisited and assessed by software architects (Falessi et al., 2013).
@@
Although the importance of DR in architecture design has been widely acknowledged, it is often inadequately recorded and detailed during software development (Tyree and Akerman, 2005; Bosch, 2004). Tang et al. (Tang et al., 2006) conducted a survey to investigate the oversight of DR by developers (e.g., “not aware of” and “no time/budget”) and the challenges of capturing DR (e.g., “no suitable tool”). Besides, extra efforts are required when architects need to capture DR (Capilla et al., 2008). The absence of DR hinders developers from understanding the reasons behind architecture decisions, thus violating the architecture decisions made in the design phase during the implementation and increased maintenance costs (Tang et al., 2007), which reveals a critical gap in the DR capture process during software architecting.
@@
The advancement of Large Language Models (LLMs) offers promising potential to address this gap by assisting in generating DR. LLMs, powered by deep learning technologies, have revolutionized Natural Language Processing (NLP) by demonstrating advanced language understanding capabilities, including syntax, semantics, and pragmatics (Brown et al., 2020). LLMs’ ability to generate diverse content stems from extensive parameters and training on large-scale datasets, enabling sustained performance improvements across a wide range of downstream tasks (Wei et al., 2022a). Previous studies (e.g., (Kabir et al., 2024; Jin et al., 2024; Widjojo and Treude, 2023)) have demonstrated LLMs’ utility in addressing Software Engineering (SE) challenges, highlighting their potential to support developers in various SE tasks. In particular, by leveraging the advanced language comprehension and reasoning capabilities of LLMs, software architects can efficiently use them to assist in generating and recovering DR of architecture decisions. Meanwhile, the accuracy and quality of DR generated by LLMs are yet to be explored. Besides, according to the study by Soliman et al. (Soliman and Keim, 2025), who employed a zero-shot method to query GPT-3.5 regarding the DR of the Hadoop Distributed File System (HDFS) [^1] https://hadoop.apache.org/docs/stable/hadoop-project-dist/hadoop-hdfs/HdfsDesign.html, GPT-3.5 exhibited moderate Recall but relatively low Precision against a predefined ground truth for DR derived from architecture issues and related documents on HDFS. Therefore, how to effectively guide LLMs to ensure the generation of high-quality DR remains a critical challenge that needs further investigation.
@@
In this paper, we conducted an empirical study to evaluate the DR of architecture decisions generated by LLMs. We first constructed a dataset containing architecture-related problems from software development scenarios. Specifically, the dataset is composed of 100 data points related to architecture decisions, including 50 posts from SO, along with 25 discussions and 25 issues from GitHub. We then chose five widely used LLMs for the evaluation, including both popular open-source and proprietary models: gpt-3.5-turbo, gpt-4-0613, gemini-1.0-pro, llama3-8B, and mistral-7B. To generate DR of architecture decisions, we employed three distinct prompting strategies: zero-shot, chain of thought (CoT), and LLM-based agents. For zero-shot and CoT methods, we designed tailored prompt templates, while for LLM-based agents, we developed a multi-agent system consisting of five LLM-based agents that collaborate to generate DR. Finally, we evaluated the DR generated by the five LLMs and the three prompting strategies using both accuracy metrics for quantitative analysis and IHUM-category for qualitative analysis.
@@
  • We constructed and publicly released a dataset containing 100 architecture-related problems from software development scenarios on SO and GitHub (Zhou et al., 2025).
@@
  • We discussed the advantages and disadvantages of the three prompting strategies, as well as the strengths and limitations of LLM-generated DR in comparison to that provided by human experts.
@@ 2.1. Definition of Design Rationale
Typically, design rationale in software development refers to the underlying reasons for making design decisions (Moran and Carroll, 2020). However, no standardized definition of DR in software design has been established, resulting in subtle variations in its interpretation across different contexts (Moran and Carroll, 2020). For example, Rogers et al. (Rogers et al., 2015) define that DR consists of the decision problem, alternative solutions, and arguments for or against the solutions. Besides, Zhao et al. (Zhao et al., 2024) defined DR as a combination of a solution and its corresponding supporting or opposing arguments.
We provided the precise interpretation of DR in this study. Our dataset is derived from the posts on SO, as well as the issues and discussions on GitHub. Each data point includes a specific architecture-related problem $\bm{P}$; an architecture decision $\bm{D}$; and its corresponding design rationale $\bm{DR}$. Design rationale ($\bm{DR}$) is composed of a set of arguments $\bm{A}$, where each argument corresponds to one particular perspective for architecture design ($\bm{D}$), such as advantages, disadvantages, and trade-offs. The relationship between architecture problem $\bm{P}$, architecture decision $\bm{D}$, and design rationale $\bm{DR}$ in the dataset is presented in Equation (1) below.
@@ 2.2. Prompt Engineering of LLMs
Prompt engineering has emerged as an essential technique for enhancing the capabilities of pre-trained LLMs (Sahoo et al., 2024). By designing prompts to guide the output, LLMs can adapt to various tasks and domains without requiring modifications to their parameters (Ratnadira et al., 2024; Deljouyi et al., 2024; Delile et al., 2023). The field of prompt engineering encompasses a wide range of techniques tailored to optimize LLM performance.
In this study, we focused on two prompting methods: zero-shot and CoT. Zero-shot is a foundational prompting technique where the task is directly presented to LLMs without requiring additional examples. This approach enables LLMs to perform tasks by leveraging its pre-training knowledge (Brown et al., 2020). We employed the zero-shot method to evaluate the performance of LLMs in generating DR using the basic prompt. On the other hand, CoT is a technique that guides LLMs through a structured sequence of reasoning steps. Specifically, the template of the CoT method instructs LLMs to generate more thoughtful and deeply reasoned answers by following a series of steps. Wei et al. (Wei et al., 2022b) demonstrated that CoT significantly improves performance in math and commonsense reasoning tasks. Given that DR often involves complex reasoning and trade-offs, we employed CoT in order to help LLMs produce more carefully considered outputs. The templates used for zero-shot and CoT methods are presented in Section 3.
@@ 2.3. LLM-based Agents
An intelligent agent is an automated entity capable of perceiving, making decisions, and taking actions (Wooldridge and Jennings, 1995). With the advent of LLMs, researchers and practitioners have increasingly focused on developing LLM-based autonomous agents, leading to the emergence of LLM-based agent frameworks (e.g., BabyAGI (Yohei, 2023) and MetaGPT (Hong et al., 2023)). These agent frameworks typically comprise multiple autonomous intelligent agents, each capable of independently performing tasks by interacting, communicating, and collaborating with other agents. Compared to prompt engineering, multi-agent systems offer more sophisticated agent interactions and coordination mechanisms. By leveraging a variety of tools, LLM-based agent systems can further enhance their capabilities, enabling the systems to collectively address more complex and dynamic challenges. Several LLM-based agent systems have been developed to tackle different tasks in SE, such as code generation (Bairi et al., 2024), code review (Hu et al., 2023), and requirements engineering (Jin et al., 2024). However, there is currently no LLM-based agent systems specifically designed for generating DR for architecture decisions. Considering the complexity of DR, we designed a multi-agent system consisting of five LLM-based agents, building on the MetaGPT framework (Hong et al., 2023). The details of this LLM-based agent system are presented in Section 3.
@@ 3. Methodology
In this section, we introduce the methodology employed in our research in terms of research questions (Section 3.1), selection of LLMs (Section 3.2), dataset construction (Section 3.3), design rationale generation (Section 3.4), and evaluation methods (Section 3.5). The overview of the research process is presented in Fig. 1.
@@ 3.1. Research Questions
Our goal is to uncover and evaluate the capability of LLMs in generating DR for architecture decisions. To this end, we formulated the following three Research Questions (RQs):
@@
Rationale: RQ1 aims to explore the feasibility and capability of using LLMs in generating DR for architecture decisions. DR justifies architecture choices and helps long-term system maintainability by providing context for software engineers to understand architecture decisions (Tang et al., 2007). LLMs can potentially automate the DR generation, helping recover and preserve the reasons behind architecture decisions.
@@
Rationale: CoT, involving a series of intermediate reasoning steps, significantly enhances the ability of LLMs to perform complex reasoning (Wei et al., 2022b). Since the CoT method enables step-by-step reasoning, it can potentially instruct LLMs to engage in more comprehensive thinking and make more refined trade-offs for architecture decisions. Therefore, we plan to explore the ability of the CoT method to enhance the quality of LLM-generated DR for architecture decisions.
@@
Rationale: LLM-based agents have been successfully applied in the field of SE, demonstrating significant effectiveness in supporting and facilitating various SE tasks (Liu et al., 2024b). By collaborating in specific roles, LLM-based agents are capable of performing complicated tasks such as reasoning, gathering information, and reviewing output. However, it remains unclear whether LLM-based agents can effectively generate DR for architecture decisions. Therefore, we designed a multi-agent system to explore whether collaboration among LLM-based agents can further enhance the ability of LLMs to generate DR for architecture decisions.
@@ 3.2. Selection of LLMs
As presented in Table 1, we selected five mainstream open-source and closed-source LLMs in this study. These LLMs exhibit high performance in natural language understanding and reasoning, and have been widely applied in many SE tasks (e.g., (Kabir et al., 2023; Kang et al., 2024; de Souza et al., 2023; Zhang et al., 2024)). Besides, the knowledge cut-off date of LLM is also an important factor in our model selection. Since SO posts, GitHub issues and discussions, are publicly available, we collected data generated after the knowledge cut-off dates of all employed LLMs to prevent data leakage issues (Balloccu et al., 2024). Note that selecting an LLM with a more recent cut-off date may result in a smaller dataset, potentially impacting the analysis of experimental results. For instance, if we choose GPT-4o, one of the most advanced models from OpenAI (OpenAI, 2024), only 28 architecture-related posts from SO, which were created after the knowledge cut-off date of GPT-4o, would remain available.
@@
Model Description Cut-off date Open-source
gpt-3.5-turbo gpt-3.5-turbo is an LLM from OpenAI’s GPT series. It can understand and generate natural language and code, excelling in tasks like conversation generation, text completion, and question answering. While optimized for chat applications, it also effectively handles non-chat tasks. 2021-09 No
gpt-4-0613 gpt-4-0613 is an advanced multimodal LLM that accepts text and image inputs and produces text outputs. It offers improved accuracy and problem-solving capabilities over previous models, leveraging its extensive knowledge and reasoning skills. gpt-4-0613 is versatile, suitable for chat, text generation, and traditional completion tasks. 2021-09 No
gemini-1.0-pro gemini-1.0-pro is an LLM developed by Google, capable of understanding and generating human language with remarkable proficiency. It can process not only text, but also understand and generate various forms of information such as images and videos. gemini-1.0-pro excels in tasks like multilingual translation, code generation, and question answering. 2023-04 No
llama3-8B llama3-8B is an open-source LLM developed by Meta for efficient natural language processing. It excels in tasks like text understanding, generation, and conversation. With a flexible design, it operates across various platforms and is freely available for researchers and developers to modify. 2023-03 Yes
mistral-7B mistral-7B is an LLM developed by Mistral AI, designed for high-performance natural language processing tasks. It excels in text generation, summarization, and understanding, offering state-of-the-art capabilities for developers and researchers. Additionally, mistral-7B is open-source, allowing for customization and integration into various projects. 2021-08 Yes

### v3

@@ Abstract.
Design Rationale (DR) for software architecture decisions refers to the reasoning underlying architectural choices, which provides valuable insights into the different phases of the architecting process throughout software development. However, in practice, DR is often inadequately documented due to a lack of motivation and effort from developers. With the recent advancements in Large Language Models (LLMs), their capabilities in text comprehension, reasoning, and generation may enable the generation and recovery of DR for architecture decisions. In this study, we evaluated the performance of LLMs in generating DR for architecture decisions. First, we collected 50 Stack Overflow (SO) posts, 25 GitHub issues, and 25 GitHub discussions related to architecture decisions to construct a dataset of 100 architecture-related problems. Then, we selected five LLMs to generate DR for the architecture decisions with three prompting strategies, including zero-shot, chain of thought (CoT), and LLM-based agents. With the DR provided by human experts as ground truth, the Precision of LLM-generated DR with the three prompting strategies ranges from 0.267 to 0.278, Recall from 0.627 to 0.715, and F1-score from 0.351 to 0.389. Besides, 64.45% to 69.42% of the arguments of DR not mentioned by human experts are also helpful, 4.12% to 4.87% of the arguments have uncertain correctness, and 1.59% to 3.24% of the arguments are potentially misleading. To further understand the trustworthiness and applicability of LLM-generated DR in practice, we conducted semi-structured interviews with six practitioners. Based on the experimental and interview results, we discussed the pros and cons of the three prompting strategies, the strengths and limitations of LLM-generated DR, and the implications for the practical use of LLM-generated DR.
@@ 1. Introduction
The software architecture of a system consists of the structures required to understand the system, including its software elements, their relationships and properties (bass2021software). When making architecture decisions, software architects need to consider multiple factors like the application domain, architectural styles and patterns, Commercial Off-The-Shelf (COTS) components, and other infrastructure choices required to meet system requirements (jansen2005software). Design Rationale (DR) explains the reasoning behind architecture decisions, encapsulating the architecture knowledge and thought processes that support the resulting design (tang2006asurvey). A comprehensive DR can support various development activities, such as change impact analysis or a major redesign. With documented DR, architecture decisions can be more easily revisited and assessed by software architects (davide2013value).
@@
Although the importance of DR in architecture design has been widely acknowledged, it is often inadequately recorded and detailed during software development (tyree2005architecture; bosch2004software). Tang et al. (tang2006asurvey) conducted a survey to investigate the oversight of DR by developers (e.g., “not aware of” and “no time/budget”) and the challenges of capturing DR (e.g., “no suitable tool”). Besides, extra efforts are required when architects need to capture DR (capilla2008effort). The absence of DR hinders developers from understanding the reasons behind architecture decisions, thus violating the architecture decisions made in the design phase during the implementation and increased maintenance costs (tang2007rationale), which reveals a critical gap in the DR capture process during software architecting.
@@
The advancement of Large Language Models (LLMs) offers promising potential to address this gap by assisting in generating DR. LLMs, powered by deep learning technologies, have revolutionized Natural Language Processing (NLP) by demonstrating advanced language understanding capabilities, including syntax, semantics, and pragmatics (brown2020language). LLMs’ ability to generate diverse content stems from extensive parameters and training on large-scale datasets, enabling sustained performance improvements across a wide range of downstream tasks (wei2022emergent). Previous studies (e.g., (kabir2024stack; jin2024mare; widjojo2023addressing)) have demonstrated LLMs’ utility in addressing Software Engineering (SE) challenges, highlighting their potential to support developers in various SE tasks. In particular, by leveraging the natural language comprehension and reasoning capabilities of LLMs (jason2022chain), software architects can efficiently use LLMs to assist in generating and recovering DR of architecture decisions. Meanwhile, the accuracy and quality of DR generated by LLMs are yet to be explored. Besides, according to the study by Soliman et al. (soliman2025do), who employed a zero-shot method to query GPT-3.5 regarding the DR of the Hadoop Distributed File System (HDFS) [^1] https://hadoop.apache.org/docs/stable/hadoop-project-dist/hadoop-hdfs/HdfsDesign.html, GPT-3.5 exhibited moderate Recall but relatively low Precision against a predefined ground truth for DR derived from architecture issues and related documents on HDFS. Therefore, how to effectively guide LLMs to ensure the generation of high-quality DR remains a critical challenge that needs further investigation.
@@
In this paper, we conducted an empirical study to evaluate the DR of architecture decisions generated by LLMs. We first constructed a dataset containing architecture-related problems from software development scenarios. Specifically, the dataset is composed of 100 data points related to architecture decisions, including 50 posts from SO, along with 25 discussions and 25 issues from GitHub. We then chose five widely used LLMs for the evaluation, including both popular open-source and proprietary models: gpt-3.5-turbo, gpt-4-0613, gemini-1.0-pro, llama3-8B, and mistral-7B. To generate DR of architecture decisions, we employed three distinct prompting strategies: zero-shot, chain of thought (CoT), and LLM-based agents. For zero-shot and CoT methods, we designed tailored prompt templates, while for LLM-based agents, we developed a multi-agent system consisting of five LLM-based agents that collaborate to generate DR. Finally, we evaluated the DR generated by the five LLMs and the three prompting strategies using both accuracy metrics for quantitative analysis and IHUM-category for qualitative analysis. Specifically, for quantitative analysis, we employed Precision, Recall, and F1-score to evaluate the generated DR, using the rationale provided by human experts as ground truth. For qualitative analysis, we introduced the IHUM-category to classify the argument points in the LLM-generated DR into four categories, allowing us to further assess the arguments not mentioned by the human experts. The detailed definition of the IHUM-category is provided in Table 4. Based on the results of our experiments, we conducted semi-structured interviews with six practitioners to understand the trustworthiness and applicability of LLM-generated DR in practice. From the perspective of practitioners, we explored their trust level in LLM-generated DR, the information that needs to be provided to LLMs to enhance their trust, their willingness to use LLM-generated DR in practice, suitable scenarios for its application, as well as the challenges and limitations associated with applying LLM-generated DR.
@@
  • We constructed and publicly released a dataset containing 100 architecture-related problems from software development scenarios on SO and GitHub (dataset).
@@
  • We conducted semi-structured interviews with practitioners to understand the trustworthiness and applicability of LLM-generated DR in practice.
  • We discussed the advantages and disadvantages of the three prompting strategies, the strengths and limitations of LLM-generated DR in comparison to that provided by human experts, as well as the implications for practical use of LLM-generated DR.
@@ 2.1. Definition of Design Rationale
Typically, design rationale in software development refers to the underlying reasons for making design decisions (moran2020design). However, no standardized definition of DR in software design has been established, resulting in subtle variations in its interpretation across different contexts (moran2020design). For example, Rogers et al. (rogers2015using) define that DR consists of the decision problem, alternative solutions, and arguments for or against the solutions. Besides, Zhao et al. (zhao2024drminer) defined DR as a combination of a solution and its corresponding supporting or opposing arguments.
We provided the precise interpretation of DR in this study. Our dataset is derived from the posts on SO, as well as the issues and discussions on GitHub. Each data point includes a specific architecture-related problem $\bm{P}$; an architecture decision $\bm{D}$; and its corresponding design rationale $\bm{DR}$. Design rationale ($\bm{DR}$) is composed of a set of arguments $\{\bm{A}_{1},\bm{A}_{2},\ldots,\bm{A}_{n}\}$, where each argument corresponds to one particular perspective for architecture design ($\bm{D}$), such as advantages, disadvantages, and trade-offs. The relationship between architecture problem $\bm{P}$, architecture decision $\bm{D}$, and design rationale $\bm{DR}$ in the dataset is presented in Equation (1). Equation (1) specifies which data items should be extracted from SO posts, GitHub issues, and GitHub discussions to construct our dataset. Accordingly, the data items D1–D3 presented in Table 3 correspond to $\bm{P}$, $\bm{D}$, and $\bm{DR}$, respectively. Moreover, Equation (1) defines DR as comprising multiple argument points. By decomposing DR into individual arguments, we can evaluate these arguments using both accuracy metrics and the IHUM-category.
@@ 2.2. Prompt Engineering of LLMs
Prompt engineering has emerged as an essential technique for enhancing the capabilities of pre-trained LLMs (sahoo2024systematic). By designing prompts to guide the output, LLMs can adapt to various tasks and domains without requiring modifications to their parameters (widyasari2024codeagent; deljouyi2024leveraging; delile2023evaluating). The field of prompt engineering encompasses a wide range of techniques tailored to optimize LLM performance.
In this study, we focused on two prompting methods: zero-shot and CoT. Zero-shot is a foundational prompting technique where the task is directly presented to LLMs without requiring additional examples. This approach enables LLMs to perform tasks by leveraging its pre-training knowledge (brown2020language). We employed the zero-shot method to evaluate the performance of LLMs in generating DR using the basic prompt. On the other hand, CoT is a technique that guides LLMs through a structured sequence of reasoning steps. Specifically, the template of the CoT method instructs LLMs to generate more thoughtful and deeply reasoned answers by following a series of steps. Wei et al. (jason2022chain) demonstrated that CoT improves performance in math and commonsense reasoning tasks. Given that DR often involves complex reasoning and trade-offs, we employed CoT in order to help LLMs produce more carefully considered outputs. The templates used for zero-shot and CoT methods are presented in Section 3.
@@ 2.3. LLM-based Agents
An intelligent agent is an automated entity capable of perceiving, making decisions, and taking actions (wooldridge1995the). With the advent of LLMs, researchers and practitioners have increasingly focused on developing LLM-based autonomous agents, leading to the emergence of LLM-based agent frameworks (e.g., BabyAGI (babyagi) and MetaGPT (hong2023metagpt)). These agent frameworks typically comprise multiple autonomous intelligent agents, each capable of independently performing tasks by interacting, communicating, and collaborating with other agents. Compared to prompt engineering, multi-agent systems offer more sophisticated agent interactions and coordination mechanisms. By leveraging a variety of tools, LLM-based agent systems can further enhance their capabilities, enabling the systems to collectively address more complex and dynamic challenges. Several LLM-based agent systems have been developed to tackle different tasks in SE, such as code generation (bairi2024codeplan), code review (hu2023large), and requirements engineering (jin2024mare). However, there is currently no LLM-based agent systems specifically designed for generating DR for architecture decisions. Considering the complexity of DR, we designed a multi-agent system consisting of five LLM-based agents, building on the MetaGPT framework (hong2023metagpt). The details of this LLM-based agent system are presented in Section 3.
@@ 3. Methodology
In this section, we introduce the methodology employed in our research in terms of research questions (Section 3.1), selection of LLMs (Section 3.2), dataset construction (Section 3.3), design rationale generation (Section 3.4), evaluation methods (Section 3.5), and interviews with practitioners (Section 3.6). The overview of the research process is presented in Fig. 1.
@@ 3.1. Research Questions
Our goal is to uncover and evaluate the capability of LLMs in generating DR for architecture decisions, as well as practitioners’ perceptions of LLM-generated DR in practice. To this end, we formulated the following four Research Questions (RQs):
@@
Rationale: RQ1 aims to explore the feasibility and capability of using LLMs in generating DR for architecture decisions. DR justifies architecture choices and helps long-term system maintainability by providing context for software engineers to understand architecture decisions (tang2007rationale). LLMs can potentially automate the DR generation, helping recover and preserve the reasons behind architecture decisions.
@@
Rationale: CoT, involving a series of intermediate reasoning steps, enhances the ability of LLMs to perform complex reasoning (jason2022chain). Since the CoT method enables step-by-step reasoning, it can potentially instruct LLMs to engage in more comprehensive thinking and make more refined trade-offs for architecture decisions. Therefore, we plan to explore the ability of the CoT method to enhance the quality of LLM-generated DR for architecture decisions.
@@
Rationale: LLM-based agents have been successfully applied in the field of SE, demonstrating significant effectiveness in supporting and facilitating various SE tasks (liu2024large). By collaborating in specific roles, LLM-based agents are capable of performing complicated tasks such as reasoning, gathering information, and reviewing output. However, it remains unclear whether LLM-based agents can effectively generate DR for architecture decisions. Therefore, we designed a multi-agent system to explore whether collaboration among LLM-based agents can further enhance the ability of LLMs to generate DR for architecture decisions.
RQ4: How do practitioners perceive LLM-generated DR of architecture decisions in practice?
Rationale: If the performance of LLMs in generating DR for architecture decisions is acceptable, their use for automatic DR generation could provide practical value for software architects. However, the perception of LLM-generated DR from the practitioners’ perspective remains unclear. To address RQ4, we conducted semi-structured interviews to investigate how practitioners evaluate the DR generated by LLMs in terms of trustworthiness and applicability.
@@ 3.2. Selection of LLMs
As presented in Table 1, we selected five mainstream open-source and closed-source LLMs in this study. These LLMs exhibit high performance in natural language understanding and reasoning, and have been widely applied in many SE tasks (e.g., (kabir2023answers; Sungmin2024evaluating; de2023lessons; zhang2024using)). Besides, the knowledge cut-off date of LLMs is also an important factor in our model selection. Since SO posts, GitHub issues and discussions, are publicly available, we collected data generated after the knowledge cut-off dates of all employed LLMs to prevent data leakage issues (balloccu2024leak). Note that selecting an LLM with a more recent cut-off date may result in a smaller dataset, potentially impacting the analysis of experimental results. For instance, if we choose GPT-4o, one of the most advanced models from OpenAI (gpt-4o), only 28 architecture-related posts from SO, which were created after the knowledge cut-off date of GPT-4o, would remain available. Therefore, after considering the trade-off between model recency and dataset size, we ultimately selected the five LLMs listed in Table 1. Among the five selected models, llama3-8B is the only LLM released in 2024, while the others were all released in 2023. The time gap between llama3-8B and the earliest model, gpt-3.5-Turbo, spans one year and one month. We selected llama3-8B due to its substantial improvements over previous versions in the LLaMA series. More importantly, its knowledge cut-off date (March 2023) is earlier than that of gemini-1.0-pro (April 2023), ensuring that it does not constrain the dataset size.
@@
Model Description Cut-off date Release date Open-source
gpt-3.5-turbo gpt-3.5-turbo is an LLM from OpenAI’s GPT series. It can understand and generate natural language and code, excelling in tasks like conversation generation, text completion, and question answering. While optimized for chat applications, it also effectively handles non-chat tasks. 2021-09 2023-03 No
gpt-4-0613 gpt-4-0613 is an advanced multimodal LLM that accepts text and image inputs and produces text outputs. It offers improved accuracy and problem-solving capabilities over previous models, leveraging its extensive knowledge and reasoning skills. gpt-4-0613 is versatile, suitable for chat, text generation, and traditional completion tasks. 2021-09 2023-06 No
gemini-1.0-pro gemini-1.0-pro is an LLM developed by Google, capable of understanding and generating human language with remarkable proficiency. It can process not only text, but also understand and generate various forms of information such as images and videos. gemini-1.0-pro excels in tasks like multilingual translation, code generation, and question answering. 2023-04 2023-12 No
llama3-8B llama3-8B is an open-source LLM developed by Meta for efficient natural language processing. It excels in tasks like text understanding, generation, and conversation. With a flexible design, it operates across various platforms and is freely available for researchers and developers to modify. 2023-03 2024-04 Yes
mistral-7B mistral-7B is an LLM developed by Mistral AI, designed for high-performance natural language processing tasks. It excels in text generation, summarization, and understanding, offering state-of-the-art capabilities for developers and researchers. Additionally, mistral-7B is open-source, allowing for customization and integration into various projects. 2021-08 2023-09 Yes